1. A researcher records the number of mistakes made during a memory skills task. He finds...

1.

A researcher records the number of mistakes made during a memory skills task. He finds that the probability that participants in this study made 0 mistakes is p = .22; made 1 mistake is p = .30; made 2 mistakes is p = .16; made 3 mistakes is p = .12; and made 4 or more mistakes isp = .25. Is this probability distribution accurate?

Yes, it distributes all possible outcomes for the random variable.

No, one of the outcomes is stated as “or more.”

No, the probability distribution does not sum to 1.0.

both B and C

2.

The following are six random outcomes for a sample space: -1, -3, -3, -2, -5, and -6. What is the probability of selecting a -3 in this example?

3.

Below is the probability distribution for random variable x. Which of the following is most probable?

x

12

13

14

p( x)

0.55

0.35

0.10

3. Below is the probability distribution for the number of times students had to take the same exam before passing it. If a student must pass the exam in two tries of less to pass the class, then what is the probability of passing the class?

x

1

2

3

4

p( x)

0.40

0.40

0.10

0.10
Answer #1

1) 0.22 + 0.3 + 0.16 + 0.12 + 0.25 = 1.05

Since the sum of individual probabilities is not equal to 1, so the probability distribution is not accurate.

Option - C) No, the probability distribution does not sum to 1.0.

2) Probability of selecting a -3 = 2/6 = 1/3

3) P(X = 12) = 0.55, it is most probable.

3) P(passing) = P(X < 2)

                       = P(X = 1) + P(X = 2)

                       = 0.40 + 0.40 = 0.80